Dividends are a distribution of a company’s earnings to stockholders. In terms of budgeted dividend

payments, dividends paid would be included on the following budget(s):

Balance Sheet Cash Budget Income Statement Budget

A) Yes Yes Yes

B) No Yes Yes

C) No Yes No

D) Yes No No

34. According to the rule of 72, if you invest $1,000 at 8% then this sum will grow to ___________ in 36

years.

A) $2,000

B) $16,000

C) $4,000

D) $32,000

E) Infinity
Answer #1

Solution 1:

Dividends are a distribution of a company’s earnings to stockholders. In terms of budgeted dividend

payments, dividends paid would be included on the "Cash budget only"

Hence option C is correct.

Solution 2:

72 / [periodic interest rate] = [number of years to double principal]

72 / 8 = 9 years

It means it will take 9 years to double the principal amount.

Therefore in 36 years accumulated amount will be = 2*2*2*2 = 16 times of original principal amount

Therefore sum will grow to = $1,000*16 = $16,000

Hence option b is correct.